OpenStudy (anonymous):

Al gives 1/2 of his marbles to Bev. Bev gives 1/3 of these to Carl. Carl gives 1/4 of these to Dani. If Dani was given four marbles, how many did Al have originally?

OpenStudy (anonymous):

96.

OpenStudy (anonymous):

96

OpenStudy (unklerhaukus):

A=2B B=3C C=4D D=4 A=2(3(4(4))) = 2x3x4x4=96

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

If Dani has 4 marbles and was given 1/4 of Carl's, then Carl must have 4*4 = 16 marbles. If Carl was given 1/3 of Bev's, then Bev must have 3*16 = 48. And if Bev was given 1/2 of Al's, then Al must have 2*48 = 96.

OpenStudy (unklerhaukus):

A=2B = 2(3C) = 2(3(4D)) =2(3(4(4)))

OpenStudy (anonymous):

so i just multply the reciprocal by 4